The equation of the circle with radius 2 and center at (-1,3).

Thanks for the big help.. Kindly put the solution..

Answer by uday1435(57)

(x – g)^2 +(y – h)^2 = r^2
So for the given center and radius
(x + 1)^2 +(y – 3)^2 = 4
x^2 + 2x + 1 + y^2 – 6y +9 = 4
x^2 + y^2 + 2x – 6y + 6 = 0 is the required equation
If you still require further clarifications please write to
udayakumar.t.r@gmail.com